Solve by substitution method.
8x+5y=29
-8x+y=25

Answer by robertb(5830)

The bottom equation gives 8x = y - 25. Substitute this into the top equation:
y - 25 + 5y = 29
6y = 54
y = 9.
==> 8x = 9 - 25
8x = -16
x = -2.